Klapperich Lab:Notebook/Lab Meeting Notes/2010/01/21

From OpenWetWare

Jump to: navigation, search
Project name Main project page
Previous entry      Next entry

21 January 2010 Lab Meeting

Mincheol Presentation.

‡ Announcements

  • The safety inspection will happen next Wed at 1PM. Lab members PLEASE pitch in with Hussam and Jane and Jessie to prepare the lab according to the items on the inspection list tomorrow from 9AM (Hussam, Jessie, Jane leading).
  • Oakridge deadline is 1 Feb 2010. Have one from QQ- Anyone else?
  • Shichu Huang will be joining us in Feb.


‡ Flu R01:Integration

  • meetings every other M 1-3pm. 705.

† HDA (Lead: Jaephil, Team:Sonali)

  • Start planning R01 for Submission on 6/5/10. MM, JD, CMK.
  • Jaephil trying out the lyophilizer with just water to save on HDA reagents. Then Sonali lyophilize HDA mixes +/- trehalose.(MM)
  • HDA at 65, 66, 67, 68, 69, 70C done with C diff ATCC DNA. Works at all except 70C. Less amplification as Temp increases (MM)
  • HDA negative water reactions all have primer dimers/non-specific band at 70bp for c diff HDA (MM)
  • Jaephil to make COP chips for LOD expt with C diff DNA. (MM)
  • 2nd HDA design - include RT, dry reagent, reduced evap loss, and parallel filling, etc
  • Paper PUBLISHED online 1/12/10!
  • "prototyping - scratch" SUBMITTED "Special Issue of the Robotics and Computer-Integrated Manufacturing Journal"


† Sample Concentration (Lead: Jane, Team: Jaephil)

  • New design sent to make Rubber mold: outline octagon problem. Revised design sent yesterday.
    • MM 040808 A/PR/8/34 flu stock with PFU = 2X10^4 available to Jane to train Cassidy in ultracentrifuging all tubes
    • of virus, pour out sucrose,pipette remaining with 20uL pipette tip and pool for plaque assay for supplying virus
  • Jane working on the cell lysate control. - have many tubes now, done
  • Set up COMSOL and StarCD, look for governing equations for simulation of evaporation. Ongoing

† SPE Column Optimization for DNA/RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)

  • Sonali put out a new RNA protocol (correct SPE batter, homemade buffers, less ethanol washing, RNA Secure) before break.
  • Protocol tested with 10ng, 1ng, 0.1ng Stratagene purified human RNA (MM) and 100 uL of 1:64 dilution ATCC MDCK sup Flu A/PR/8/34 (JC)
  • Human Total RNA see % yields at 50% and fraction 1 + fraction 2 = 4-5ng RNA out of 10ng with all changes. With high vol *ethanol wash but other protocol improvements, see 40% yield at all other concentrations.
  • For Flu virus see improved Ct value by 7 cycles in fraction 1 from multiple channels
  • Also detecting 1.4 logs more of virus at 1:64 JC virus prep than before.
  • Switching from lambda phage DNA to TaqMan RNase P detection [update]
  • Hussam Runs chip with RNA (confirm procedure is correct) [update]
  • Jessie will run chips with serial virus dilutions from 1:1 to 1:1,000,000 to re-do previous dilution experimnent with WHO SYBR green assay on chip.
  • Sonali will assay BIDMC virus sample with newest protocol with 0.7uM Silica and compare with Qiagen RNA extraction with CDC PCR assay


† PCR - CMI (Lead: Qingqing)

  • 6th Integrated chip with SPE+RT+PCR chip was designed.
  • "crack" issue is solved.
  • It works for ATCC flu virus, however low efficiency.
  • working on improving the efficiency

† PCR of C.Difficile DNA (Qingqing)

  • Meeting today at 3:30pm
  • Toxin A.
  • HDA dilution experiments for toxin A has been tested in tube. The products has been tested in Bio-analyzer. 1pg is the limitation for the template, in order to get a detectable result.
  • Toxin B.
  • PCR assay works on chip with ATCC C.dif DNA.
  • old SPE DNA has ct values above 30, on chip PCR with these DNA does not work either on 30 cycle chip nor on 40 cycle chip.
  • Qiagen extracted DNA from Lisa on Jan 6 have ct values between 20-30.
  • on chip PCR with lisa's DNA also does not work either on 30 cycle chip nor on 40 cycle chip.
  • PCR2 Paper back to MCK this week.
  • CMK: PCR 1 draft. Analytical Chem.


‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )

  • See email notes.


‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)

  • New meeting time, T, 9:30 am every other week.


‡ Agilent Automated Sample Preparation (Lead: Alex)

  • Paper on HotDog in progress. First Draft by 1/24
  • NANO-C:
    • In Progress: More tests with B. Subtillis until 1/20


‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)

  • New metal piece - First week of Dec.
  • Evap paper -CMK edited, back to JD and JZ this week


‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
* Cathie will submit paper inquiry.

‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)

  • First samples collected.


‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))

  • Submitted Gantt Document to PATH 11/25.
  • Frank: 1-week in-tube stabilization experiment successful. Met with Sonali to learn hood cleaning and verify straw extraction protocol. Doing extractions for 1-week storage today (21 Jan).
  • Sean: Ready to start preparing the monoliths today (1.21) after the meeting. Will be able to make it to the meeting by 3 every week. Hussam will be training me on PCR next week. Need bench training still
  • Mark to begin initial DNA testing with shorter time periods (1,3 days, etc.) to be completed this(?) week. New straw fixture to be set up in old fume hood. Fume hood to be sterilized for use only with straw machine.


‡ IIH Senior Project.

  • Ordering ELISA Kits
  • Determining materials to test fluorescence
    • Transparancy, COP, PDMS (gold-standard)
    • material list from study (Xurography, Daniel A. Bartholomeusz, p 1367)
  • Planned to reconnect with Jose, he is out of town until feb.
  • Returning to the cutter plotter at MIT for 'origami testing'
    • microscope?




Personal tools